# Adaptive Multi-modal Fusion of Spatially Variant Kernel Refinement with Diffusion Model for Blind Image Super-Resolution

### v1

@@ 4.1 Experiment Setup
Implementation details. In accordance with the prevailing blind super-resolution methodology, DKENet employed a $21\times 21$ anisotropic Gaussian blur kernel in the experimental setup. Within the SVKR, $T$ was set as 3, and $\gamma$ was set to 0.9. For monocular depth estimation, we use code from [42, 4]. In the context of the AMF, the hidden layer dimensions of $M_{\alpha}$ and $M_{\beta}$ were specified as 128, featuring a ReLU activation layer. AMF was incorporated into the different levels of the UNet within the Diffusion Model.
@@
Figure 7: Visual comparisons of several representative methods on examples of the blackberry dataset. Zoom in for best view.
Evaluation Metrics. The PSNR and SSIM metrics are widely employed in the field of image restoration to gauge the degree of image similarity. Thus, we employ these metrics to assess the restoration abilities of our method for high-resolution images. However, it should be noted that in the realm of human perception and considering a multitude of distortions, they may fall short of faithfully representing image quality. Hence, we additionally utilize the LPIPS [75], CLIP-IQA [51], NIQE [38] and MUSIQ [21] metrics to appraise the perceptual quality of the super-resolved images.
@@ 4.2 Comparison With Existing Methods
Compared Methods. In order to validate the performance and efficacy of our methodology, we conducted comparative analyses of SSR against the contemporary state-of-the-art and representative blind super-resolution methods, which encompass DASR [52], DCLS [33], Real-ESRGAN [54], MANet [28], SwinIR [27], HAT [9], DAN [31], DDNM [61] and ResShift [69]. For these methodologies, we executed assessments employing their official source code. Among these, DAN and DCLS represent prominent kernel estimation techniques, MANet is a pioneering venture into spatially variant blur kernels, and DASR engages in super-resolution through the utilization of abstract degradation information. ResShift and DDNM, on the other hand, is a diffusion-based model.
Quantitative Comparisons. As Table 2 demonstrates, the performance of SSR is outstanding across multiple datasets. The experimental data reveal that SSR consistently surpasses other methods in the DIV2K, BSDS100, and Urban100 datasets. Specifically, within the DIV2K dataset, SSR achieved the highest PSNR at 22.77 dB, illustrating its exceptional capability in image quality reconstruction. Likewise, SSR secured the top spot in SSIM with a score of 0.75, signifying its precision in maintaining image structure. With an LPIPS score of 0.30, SSR significantly outperforms its competitors, with a lower score indicating superior perceived image quality. SSR also leads in the no-reference image quality assessment indices of MUSIQ, NIQE, and CLIP-IQA. Moreover, on the BSDS100 and Urban100 datasets, SSR dominates in the majority of metrics, affirming its perceptual superiority. More Quantitative comparision results will be offered in the supplementary materials.
Qualitative Comparisons. Figure 5, 6 and 7 present visual comparison of the SSR methodology against established image super-resolution techniques such as RealESRGAN, DCLS, and SwinIR. Within the experimental framework, SSR exhibits superior fidelity and an impressive capacity for detail retention, particularly in the preservation of edges and intricate textures. Previous methodologies, notably DCLS and MANet, demonstrate high sensitivity to noise within low-resolution images, resulting in suboptimal visual outcomes—a deficiency aptly reflected in metrics like MUSIQ. SSR consistently outperforms other methods in these respects. The images generated by SSR exhibit notably higher clarity and sharpness, indicating its effectiveness in minimizing artifacts and distortions commonly introduced during the super-resolution process. For instance, the first row of Figure 5, illustrates how conventional methods grapple with blurring (DCLS, MANet), undue smoothness (Real-ESRGAN), and noise (ResShift) in the reconstruction of the eye and surrounding hair textures. SSR demonstrates a commendable balance between sharpness and naturalness, avoiding overly smoothed or exaggerated textures. More Qualitative comparision results will be offered in the supplementary materials.
@@ 4.3 Ablation Study
Importance of the SVKR. Our investigation initially focused on the impact of Depth-Informed Kernel (DI-Kernel), Depth Map, and SVKR on the ultimate super-resolution outcomes. As demonstrated in Table 2, when solely integrating either DI-Kernel or Depth Map, a significant deterioration in SSR performance was observed. However, when amalgamating information from all three modalities, the SSR’s PSNR and SSIM reached 21.98 dB and 0.71, respectively. The incorporation of SVKR resulted in enhancements of 3.59% in PSNR and 6.20% in SSIM for the SSR. This indicates that in SSR, multimodal features synergistically supplement missing information, and SVKR effectively enhances the accuracy of information across different modalities.
@@
Importance of the Depth-Inforemed Kernel. Subsequently, we examined the significance of accurately estimating the DI-Kernel. In Figure 8, SSR employs a blur kernel $K$ estimated by DKENet for image super-resolution, where (a) and (b) represent super-resolution using $K_{1}$ and $K_{2}$ respectively. $K_{1}$ and $K_{2}$ denote deviations from $K$ in negative and positive directions, symbolizing inaccurate blur kernels. As illustrated in Figure 8, when the blur kernel estimation is imprecise, there is a marked degradation in both visual quality and evaluation metrics of the image.
Figure 8: Ablation Comparison Between Using Accurate and Deviated Blur Kernels. (a) represents super-resolution using $K_{1}$, (b) represents super-resolution using $K_{2}$, and SSR denotes super-resolution using $K$ estimated by DKENet. $K_{1}$ and $K_{2}$ respectively signify deviations from $K$ in different directions. In visualization, we calculate the variance of the blur kernel to illustrate its intensity variations and then present this data in a heatmap.
@@
Figure 9: Ablation Comparison Between Using Accurate and Deviated Depth Map. (a) represents super-resolution using $D_{1}$, (b) represents super-resolution using $D_{2}$, and SSR denotes super-resolution using $D$ estimated by Monocular depth estimate algorithm. $D_{1}$ and $D_{2}$ respectively signify deviations from $D$ in different directions.
Performance under various degradation modes. To validate the adaptability of SSR to image super-resolution tasks under varying degradation modes, we conducted the experiments presented in Table 3. These results showcase the adaptability of the image super-resolution methodology under different degradation conditions. In scenarios devoid of noise and JPEG compression, the PSNR and SSIM metrics exhibited similar performances, demonstrating the method’s consistent efficacy across diverse conditions. Overall, these outcomes affirm that the super-resolution method can effectively adapt to a multitude of degradation modes, positively impacting image quality in practical applications.
Table 3: Quantitative results of SSR on the DIV2K dataset under different degradation models.
@@
Table 4: Comparison on DPED-iphone. Bold indicates the best performance
Method Backbone PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ CLIP-IQA $\uparrow$ NIQE $\downarrow$ Time
RealESRGAN GAN 25.24 0.82 0.33 60.10 0.41 3.28 5s
LDM Diffusion Model 24.35 0.77 0.36 52.77 0.29 3.57 55s
SR3 Diffusion Model 9.39 0.23 0.65 29.96 0.31 13.30 50s
DDNM Diffusion Model 19.34 0.64 0.56 17.03 0.17 10.54 613s
SSR Diffusion Model 25.01 0.86 0.31 62.58 0.55 3.07 32s
@@ 5 Conclusion and Discussion
The SSR framework represents a significant advancement in the field of image super-resolution. By integrating depth information and addressing the spatially variant in blur kernels, this framework substantially improves the authenticity of super-resolution images. The novel components, such as the Depth-Informed Kernel, the Spatially Variant Kernel Refinement (SVKR), and the Adaptive Multi-Modal Fusion (AMF) module, work in concert to enhance the precision and realism of super-resolution outcomes. The empirical evidence from quantitative and qualitative analyses, along with ablation studies, underscores the efficacy and superiority of our approach, marking a notable contribution to the domain of image processing and super-resolution technology.
As shown in Table 4, although the SSR achieves remarkable performance, it is limited by the diffusion model; SSR requires 30 seconds to generate a $768\times 768$ image. Model compression techniques such as distillation and pruning are promising directions to address these issues. On the other hand, SSR also holds developmental potential in other low-level vision tasks, such as deblurring and de-jittering, which are reserved for our future work.

### v2

@@ 4.1 Experiment Setup
Implementation details. In accordance with the prevailing blind SR methodology, DKENet employed a $21\times 21$ anisotropic Gaussian blur kernel in the experimental setup. Within the SVKR, $T$ was set as 3, and $\gamma$ was set to 0.9. For monocular depth estimation, we use code from [43, 4]. In the context of the AMF, the hidden layer dimensions of $M_{\alpha}$ and $M_{\beta}$ were specified as 128, featuring a ReLU activation layer. AMF was incorporated into the different levels of the UNet within the Diffusion Model.
@@
Figure 7: Visual comparisons of several representative methods on examples of the DPED-blackberry dataset. Zoom in for best view.
Evaluation Metrics. The PSNR and SSIM metrics are widely employed in the field of image restoration to gauge the degree of image similarity. However, it should be noted that in the realm of human perception and considering a multitude of distortions, they may fall short of faithfully representing image quality. Hence, we additionally utilize the LPIPS [69], CLIP-IQA [52], NIQE [39] and MUSIQ [20] metrics to appraise the perceptual quality of the images.
@@ 4.2 Comparison With Existing Methods
Quantitative Comparisons. As Table 2 demonstrates, the performance of SSR is outstanding across multiple datasets. The experimental data reveal that SSR consistently surpasses other methods in the DIV2K, BSDS100, and Urban100 datasets. Specifically, within the DIV2K dataset, SSR achieved the highest PSNR at 22.77 dB, illustrating its exceptional capability in image quality reconstruction. Likewise, SSR secured the top spot in SSIM with a score of 0.75, signifying its precision in maintaining image structure. With an LPIPS score of 0.30, SSR significantly outperforms its competitors, with a lower score indicating superior perceived image quality. SSR also leads in the no-reference image quality assessment indices of MUSIQ, NIQE, and CLIP-IQA. Moreover, on the BSDS100 and Urban100 datasets, SSR dominates in the majority of metrics, affirming its perceptual superiority.
Qualitative Comparisons. Figure 5, 6 and 7 present visual comparison of the SSR methodology against established image SR techniques such as RealESRGAN, DCLS, and SwinIR. Within the experimental framework, SSR exhibits superior fidelity and an impressive capacity for detail retention, particularly in the preservation of edges and intricate textures. Previous methodologies, notably DCLS and MANet, demonstrate high sensitivity to noise within LR images, resulting in suboptimal visual outcomes—a deficiency aptly reflected in metrics like MUSIQ. SSR consistently outperforms other methods in these respects. The images generated by SSR exhibit notably higher clarity and sharpness, indicating its effectiveness in minimizing artifacts and distortions commonly introduced during the SR process. For instance, the first row of Figure 5, illustrates how conventional methods grapple with blurring (DCLS, MANet), undue smoothness (Real-ESRGAN), and noise (ResShift) in the reconstruction of the eye and surrounding hair textures. SSR demonstrates a commendable balance between sharpness and naturalness, avoiding overly smoothed or exaggerated textures.
@@ 4.3 Ablation Study
Importance of the SVKR. Our investigation initially focused on the impact of Depth-Informed Kernel (DI-Kernel), Depth Map, and SVKR on the ultimate SR outcomes. As demonstrated in Table 2, when solely integrating either DI-Kernel or Depth Map, a significant deterioration in SSR performance was observed. However, when amalgamating information from all three modalities, the SSR’s PSNR and SSIM reached 21.98 dB and 0.71, respectively. The incorporation of SVKR resulted in enhancements of 3.59% in PSNR and 6.20% in SSIM for the SSR. This indicates that in SSR, multimodal features synergistically supplement missing information, and SVKR effectively enhances the accuracy of information across different modalities.
@@
Importance of the Depth-Inforemed Kernel. Subsequently, we examined the significance of accurately estimating the DI-Kernel. In Figure 8, SSR employs a blur kernel $K$ estimated by DKENet for image SR, where (a) and (b) represent SR using $K_{1}$ and $K_{2}$ respectively. $K_{1}$ and $K_{2}$ denote deviations from $K$ in negative and positive directions, symbolizing inaccurate blur kernels. As illustrated in Figure 8, when the blur kernel estimation is imprecise, there is a marked degradation in both visual quality and evaluation metrics of the image.
Figure 8: Ablation Comparison Between Using Accurate and Deviated Blur Kernels. (a) represents SR using $K_{1}$, (b) represents SR using $K_{2}$, and SSR denotes SR using $K$ estimated by DKENet. $K_{1}$ and $K_{2}$ respectively signify deviations from $K$ in different directions. In visualization, we calculate the variance of the blur kernel to illustrate its intensity variations and then present this data in a heatmap.
@@
Figure 9: Ablation Comparison Between Using Accurate and Deviated Depth Map. (a) represents SR using $D_{1}$, (b) represents SR using $D_{2}$, and SSR denotes SR using $D$ estimated by Monocular depth estimate algorithm. $D_{1}$ and $D_{2}$ respectively signify deviations from $D$ in different directions.
Performance under various degradation modes. To validate the adaptability of SSR to image SR tasks under varying degradation modes, we conducted the experiments presented in Table 3. These results showcase the adaptability of the image SR methodology under different degradation conditions. In scenarios devoid of noise and JPEG compression, the PSNR and SSIM metrics exhibited similar performances, demonstrating the method’s consistent efficacy across diverse conditions. Overall, these outcomes affirm that the SR method can effectively adapt to a multitude of degradation modes.
Table 3: Results of SSR on the DIV2K dataset under different degradation.
@@ 5 Conclusion and Discussion
The SSR framework represents a significant advancement in the field of image SR. By integrating depth information and addressing the spatially variant in blur kernels, this framework substantially improves the authenticity of SR images. The empirical evidence from quantitative and qualitative analyses, along with ablation studies, underscores the efficacy and superiority of our approach, marking a notable contribution to the domain of image processing and SR technology. SSR also holds developmental potential in other low-level vision tasks, such as deblurring and de-jittering, which are reserved for our future work.
Acknowledgements
This work was supported by National Natural Science Foundation of China (No.62072112), Scientific and Technological innovation action plan of Shanghai Science and Technology Committee (No.22511102202), National Key Research and Development Program of China (2023YFC3604802), China Postdoctoral Science Foundation under Grant (2023M730647, 2023TQ0075).